A and B are events with Pr[AB] = 0.5, PrIAB] = 0.3, PrIABI = 0.1, and Pr[ABl = 0.1. a. What are Pr[Al and Pr[B]? b. What are PrA B PrlAB], Pr[A|B], and PrAB? c. Calculate Pr[B|A] from Pr[A B using Bayes theorem. 2.11
Answer #1

a)

P( A) = P(AB) P(AB) = 0.3 + 0.5 = 0.8

and, P(B) = P(AB) + P(AB) = 0.1 + 0.5 = 0.6

b)

PAB)PAB) 0.5 PAB) 0.5 9222 (B) 0,6 0.8333

P(AB0.3 PlAB)--P(B) = 1-0.6 0.75

P(AB) 0.1

P(B) 1-0.6-0.25

c)

P(BA)P(B).PAB 0625 P(B).P(AB) P( B) PA B) + P(万).PlA B) 0.6(0.8333) O 6(0.8333) + (1-0.6)0.75

P(A).P(BA) P(A).PlBIA)+P(A).PlBlA) 0.8(0.625) 0.8(0.625) +0.1 P(AB)